# FuseFormer: A Transformer for Visual and Thermal Image Fusion

### v1

@@ IV Results
Attention Mech Self Axial N/A%
In this paper, also model tuning experiments are made and the results are given in Table III. For model tuning, affect of changing learning rate, batch size, transformer layers and attention mechanisim are investigated. According to Table IV, changing the number of layers on Transformer side from 8 to 12 improves the resulting SSIM. Due to nature of Transformers, this improvements trades off a huge computation resource and increases the convergence time quadratic. According to Table V, as expected, changing the batch size does not affect the model performance much. Considering the Transformer’s huge computation need, smaller batch sizes comes with a trade off of increased training time and smaller computation power. According to Table VI, changing the learning rate, especially adaptively changing during the training, improves the quality of resulting model. While starting with larger learning rate value, decreasing the learning rate towards to the end improves the convergence.
TABLE IV: Model Tuning Results: Transformer Layers
Experiment Entropy[42]normal-$\uparrow$ SCD[43]normal-$\downarrow$ MI[44]normal-$\uparrow$ SSIM[28]normal-$\uparrow$
$T_{12}$ 4.533 6.152 1.346 0.846
$T_{10}$ 4.535 6.335 1.232 0.824
$T_{8}$ 4.587 6.657 0.874 0.719
TABLE V: Model Tuning Results: Batch Size
Experiment Entropy[42]normal-$\uparrow$ SCD[43]normal-$\downarrow$ MI[44]normal-$\uparrow$ SSIM[28]normal-$\uparrow$
$B_{4}$ 4.542 6.478 1.130 0.746
$B_{2}$ 4.587 6.657 0.874 0.719
TABLE VI: Model Tuning Results: Learning Rate
Experiment Entropy[42]normal-$\uparrow$ SCD[43]normal-$\downarrow$ MI[44]normal-$\uparrow$ SSIM[28]normal-$\uparrow$
$LR_{1e-6}$ 4.553 6.551 1.032 0.774
$LR_{1e-5}$ 4.570 6.603 0.948 0.747
$LR_{1e-4}$ 4.587 6.657 0.874 0.719
@@ V Conclusion
In this paper, we proposed FuseFormer, a transformer based visible-infrared band image fusion network, where we developed a novel fusion strategy with a novel loss function that contributes from both input images and takes global context into account. In the dual-branch fusion strategy, a CNN and a transformer branches are introduced to fuse both local and global features. The proposed method is evaluated on multiple fusion benchmark datasets where we achieve competitive results compared to the existing fusion methods.

### v2

@@ IV Experiments and Results
IV-A Model Tuning Experiments
In the last round of our experiments, we investigate the impact of adjusting certain hyperparameters to fine-tune our model for improved functionality. First, we provide the quantitative performance of the models with 8, 10, and 12 transformer layers in Table III to assess the effect of increasing the number of transformer layers. It is seen that as the capacity of the model is increased there is still room for improvement. However, this improvement comes with a huge computation demand and increases the convergence time quadratically. Table IV shows that, changing the batch size does not affect the model performance significantly, which is expected. According to Table V, changing the learning rate adaptively during training, improves quantitative results. Our experience with the model also demonstrates that convergence is improved by beginning with a greater learning rate value and gradually lowering it as training progresses. We finally describe the improvements achieved by adjusting these three hyperparameters in Table VI. A key takeaway from the tuning trials is that more processing power for training can enhance the quantitative results even more.
(a) Visual-band Images
(b) Infrared Images
(c) FuseFormer
(d) SwinFusion[27]
(e) M3FD[51]
(f) RFN-Nest[44]
(g) DenseFuse[8]
Figure 6: Visual Comparison of the SoTa methods for general conditions. The two right-most scenarios present low-light conditions.
@@ V Conclusion
In this paper, we proposed the “FuseFormer” architecture, which is a transformer+CNN-based visible-infrared band image fusion network. Our proposed dual-branch fusion strategy utilizes a CNN and a transformer branch, which fuse local and global features. The proposed method is evaluated on fusion benchmark datasets where we achieve competitive results compared to the existing fusion methods. We also develop a novel fusion strategy with a novel loss function that contributes from both input (i.e. visible-band and infrared) images and takes global context into account. We even achieve better quantitative results using the original SSIM metric, despite updating this metric in the loss function to benefit from both the visible-band and the infrared image.
Image fusion presents a fundamental challenge directed towards accommodating human observers rather than algorithms, requiring additional investigation across diverse spectral bands and scenarios. However, the fusion process can also be used to improve downstream activities such as feature fusion. With the increasing integration of vision transformers, we believe that there’s optimism for achieving a universal fusion solution independent of the input bands and scenarios.